# Interstellar medium gas heating by primordial black holes and dark matter particles

### v1

@@ II.4 Gas heating by heavy PBHs
where $n_{\rm H}r_{\rm sys}=7.56\times 10^{19}$ cm-2 represents the gas column density in Leo T. The heating rate is estimated as
@@
where $f_{h}$ is taken to be $1/3$. The kinetic energy of a proton ($m_{p}=1.67\times 10^{-27}$ kg = 938.2 MeV) at a radius $r$ is determined based on the Keplerian velocity as
$E(r)\simeq\frac{1}{2}m_{p}v^{2}=782~{}{\rm keV}\left(\frac{r}{3r_{s}}\right)^{-1}\left(\frac{f_{k}}{0.1}\right)^{2},$ (15)
where $r_{\rm s}=2GM$ is the Schwarzschild radius. We approximate the Keplerian velocity as $v(r)\simeq f_{k}\sqrt{GM/r}$ with $f_{k}=0.2$. The inner disk radius is taken to be the innermost stable circular orbit (ISCO) of a test particle, $r_{\rm in}=r_{\rm ISCO}=3r_{\rm s}$. The outer radius is set to $r_{\rm out}=100r_{s}$. The energy distribution function of the outflow is given by
@@
×$\times\left(\frac{n_{\rm H}}{1~{}{\rm cm^{-3}}}\right)\left(\frac{\tilde{v}}{10~{}{\rm km/s}}\right)^{-3}\left(\frac{r_{\rm out}}{r_{s}}\right)^{-s}f_{v}^{2s}E^{-s-1},$ (16)
where the duty cycle factor is set to 1, $\tilde{v}=(v^{2}+c_{s}^{2})^{1/2}\simeq 10$ km/s in Leo T (with $v$ and $c_{s}\simeq 9$ km/s representing the PBH velocity relative to the gas and the temperature-dependent sound speed in the gas, respectively), and $f_{v}$ denotes the velocity distribution of the PBHs. Given the lack of significant relative bulk velocity in Leo T, we assume a standard Maxwell-Boltzmann distribution in polar velocity coordinates as
@@
where $\sigma_{v}$ is the velocity dispersion. Since the velocity dispersion of the PBHs is expected to match that of the gas, we take $\sigma_{v}=\sigma_{g}=6.9$ km/s, giving an average velocity $\bar{v}=\sigma_{v}\sqrt{8/\pi}\simeq 11.0$ km/s. The parameter $s$ has been estimated by simulations to range from 0.4 - 0.8. In this study, we use $s=0.7$.
@@
It is important to note that heavy PBHs cannot constitute all of the DM. The maximum fraction of heavy PBH to DM is $f_{\rm PBH}\simeq 0.08$, which is only allowed within the narrow mass range of $1\leq M_{\rm PBH}/M_{\odot}\leq 30$ in green band in FIG. 2. In other mass regions, heavy PBHs can only make up a very small fraction of the DM Carr2021RPP ; Takhistov2022JCAP . Therefore, similar to the case with light PBHs, the actual heating ratio in Leo T must be adjusted downward from what is shown in FIG. 2. However, unlike the light PBHs scenario, the case of heavy PBHs may yield significant results for the remainder of this paper, because $Q_{\rm PBH,outflow}$ can be comparable to $C$ in this mass range.
@@
II.5 Gas heating by dark photons
@@
In this subsection, we consider dark photons as the sole heating source for the gas in Leo T. Dark photons $A^{\prime}$ are kinematically mixed with the standard model photon $A$ Dubovsky2015JCAP . The kinetic mixing results in a direct coupling between the dark photon and the standard model electric current $J^{\mu}_{\rm em}$. According to Ref.Dubovsky2015JCAP ; Wadekar2021PRD , the relevant Lagrangian for this study is
$\mathcal{L}=$ $-\frac{1}{4}{F}^{\mu\nu}F_{\mu\nu}-\frac{1}{4}{F^{\prime}}^{\mu\nu}F^{\prime}_{\mu\nu}+\frac{m_{A^{\prime}}^{2}}{2}A^{\prime\mu}A^{\prime}_{\mu}$
$-\frac{e}{(1+g^{2})^{1/2}}J^{\mu}_{\rm em}\left(A_{\mu}+gA^{\prime}_{\mu}\right),$ (19)
where $m_{A^{\prime}}$ denotes the mass of the dark photon, $g$ is the coupling parameter that describes the strength of the kinetic mixing, and $A_{\mu}$ ($A^{\prime}_{\mu}$) and $F_{\mu\nu}$ ($F^{\prime}_{\mu\nu}$) represent the gauge fields and field strengths of the visible (dark) photons, respectively.
@@
where $n_{e}$ denote the number density of electron, $T_{e}$ is the electron temperature, $\alpha_{\rm em}=1/137$ is the fine structure constant, and $\ln\Lambda_{\rm C}=0.5\ln[4\pi T_{e}^{3}/(\alpha_{\rm em}^{3}n_{e})]$ is the Coulomb logarithm, respectively.
@@
II.6 Gas heating by millicharged particles
In this subsection, we consider millicharged particles as the sole heating source for the gas in Leo T. In the millicharged DM model, a millicharged particle $\chi$ possesses a small electric charge $\epsilon e$ where $e$ is the electron charge Holdom1986PLB . There are two primary ways in which millicharged particles interact with the gas in Leo T Wadekar2021PRD .
The first interatcion is Coulomb-like, involving the millicharged particles and free electrons or ions. In the gas of Leo T, almost all ions are ionized hydrogen atoms ${\rm H^{+}}$ (proton $p$). The differential Rutherford scattering cross section for a millicharged particle $\chi$ interacting with an electron or ionized hydrogen is given by Wadekar2021PRD
@@ II.6 Gas heating by millicharged particles
$\lambda_{D}=\sqrt{\frac{T}{4\pi\alpha_{\rm em}\sum_{i}Z_{i}^{2}n_{i}}}.$ (24)
The second interaction occurs when the de Broglie wavelength of millicharged particle becomes smaller than the atom’s screening length. In this regime, the millicharged particle can interact with the nucleus of neutral atoms, as the nucleus is screened over a distance approximately equal to the Bohr radius $a_{0}$. For hydrogen atoms, this interaction becomes significant when $m_{\chi}\gtrsim 0.07$ GeV. In Leo T, neutral hydrogen and helium atoms are more prevalent than electrons or ions. Thus, for $m_{\chi}\gtrsim 0.07$ GeV, the dominant heat exchange by milicharged particles comes from neutral H and He atoms Wadekar2021PRD . According to Ref.Kouvris2014PRD ; Wadekar2021PRD , the differential cross section for a millicharged particle $\chi$ interacting with hydrogen atom or helium atom is given by
$\sigma_{i\chi}=$ $\frac{2\pi Z_{i}^{2}\alpha_{\rm em}^{2}\epsilon^{2}}{\mu_{i}^{2}v_{\rm rel}^{4}}\left[\ln(1+4\mu_{i}^{2}v_{\rm rel}^{2}a^{2})-\frac{1}{1+(4\mu_{i}^{2}v_{\rm rel}^{2}a^{2})^{-1}}\right]$
${\rm for}\ i=({\rm H,He}),$ (25)
@@
We assume that both the millicharged particle $\chi$ and the gas component $i$ follow Maxwell-Boltzmann distributions in Leo T. Under this assumption, the millicharged particle can heat (or cool) gas if $T_{i}\neq T_{\chi}$. Since there is no relative bulk velocity between the millicharged particle and gas in Leo T, the heating rate of the gas component $i$ per unit volume per unit time can be expressed as Wadekar2021PRD ; Dvorkin2014PRD
@@
where we consider the following form for the cross section of interactions between the millicharged particle and the gas component $i$ interactions
$\sigma_{i\chi}=\sigma_{i\chi}^{(0)}v_{\rm rel}^{n},$ (27)
with $n=-4$. Since the rms line-of-sight velocity dispersions of stars and the ${\rm H_{I}}$ gas are similar in Leo T, the velocity dispersions of the gas and the millicharged particle are approximately equal $v_{\chi}\simeq v_{\rm H_{I}}$. Consequently, thermal exchange between the millicharged particle and the gas depends on the mass differences because the millicharged particle and the proton, since $T_{\chi}-T_{\rm H_{I}}\simeq(m_{\chi}-m_{p})v_{\rm H_{I}}^{2}$. Thus, the heating rate will be small for $m_{\chi}\simeq m_{p}\simeq 1$ GeV.
@@
FIG. 4 illustrates this heating rate, assuming all DMs is composed of millicharged particles. The curve represents the heating rate under this assumption, while the horizontal line indicates the gas cooling rate.
III Gas heating by both PBHs and DM particles
III.1 PBHs and dark photons
@@
The top panel of FIG. 5 illustrates the constraints on the PBH fraction $f_{\rm PBH}$ imposed by interstellar medium gas heating from both PBHs and dark photons in the Leo T dwarf galaxy. The black curve represents the upper limit of $f_{\rm PBH}$ when ignoring the gas heating effect of dark photons, as reported in Ref.Takhistov2022JCAP . The remaining three curves show the upper limit of $f_{\rm PBH}$ considering gas heating from both PBHs and dark photons. The red, blue, and green curves correspond to dark photon masses $m_{A^{\prime}}$ of $7\times 10^{-15}$ eV, $7.1\times 10^{-15}$ eV, and $7.2\times 10^{-15}$ eV, respectively. The excluded regions for $f_{\rm PBH}$ are indicated from gravitational waves LIGOVIRGO2019PRL (GW : red), Icarus Oguri2018PRD (I : yellow), combined bounds from surviving astrophysical systems in Eridanus II Brandt2016APJL , Segue 1 Koushiappas2017PRL , and disruption of wide binaries Rodriguez2014APJ (S : orange), and Plank Haimoud2017PRD ; Serpico2020PRR (P : gray).
@@
Next, we investigate the correlation between the upper bound on the mixing parameter $g$ and the mass of the dark photon $m_{A^{\prime}}$. We focus on the scenario where both PBHs and dark photons contribute effectively to the gas heating in Leo T. We consider PBH masses in the range $1\leq M_{\rm PBH}/M_{\odot}\leq 30$. In this mass range, the gas heating rate may be comparable to the cooling rate $C\simeq Q_{\rm PBH,outflow}$, see FIG. 2. Additionally, we use $f_{\rm PBH}=0.08$ as a benchmark case Carr2021RPP ; Takhistov2022JCAP .
The bottom panel of FIG. 5 shows the constraints on the kinetic mixing parameter $g$ from interstellar medium gas heating by PBHs and dark photons in the Leo T dwarf galaxy. The black curve represents the upper limit on $g$ when ignoring the gas heating effects of dark photons, as previously discussed in Ref. Wadekar2021PRD . The remaining three curves depict the upper limit on $g$ considering gas heating by both PBHs and dark photons. The red, blue, and green curves correspond to PBH masses of $10M_{\odot}$, $20M_{\odot}$, and $30M_{\odot}$, respectively. Excluded regions for $g$ are also shown, including constraints from the Milky Way interstellar medium Dubovsky2015JCAP (MW : orange), the gas cloud G33.4-8.0 Wadekar2021PRD (G : yellow), and the CMB Arias2012JCAP ; Jaeckel2010ARNPS (CMB : gray).
We observe that most significant constraint on the kinetic mixing parameter is already obtained even when ignoring the gas heating effect of PBHs Wadekar2021PRD . When considering the gas heating effect of PBHs in Leo T, this constraint is modified but remains the most significant.
III.2 PBHs and millicharged particles

### v2

@@ 2.4 Gas heating by heavy PBHs
where $n_{\rm H}r_{\rm sys}=7.56\times 10^{19}$ cm-2 represents the gas column density in Leo T.
The outflows reduce the falling gases and the accretion rate at smaller radii. The outflow rate can be approximated by a self-similar power-law [62]
$\dot{M}_{\rm out}(r)=\dot{M}_{\rm in}(r_{\rm out})\left(\frac{r}{r_{\rm out}}\right)^{s},$ (14)
where $r_{\rm out}$ is outer radius and $\dot{M}_{\rm in}(r_{\rm out})$ is the Bondi-Hoyle-Lyttleton accretion rate. The energy deposit from streaming outflow of proton can be estimated by the convolution of the proton emission with the heat generated per proton $\Delta E$. According to the formalism in Ref. [25, 26], the heating rate is estimated as
@@
where, $\mathcal{D}$ denotes the duty cycle factor, $\mu$ denotes the mean molecular weight. The fraction of energy deposited as heat $f_{h}$ is taken to be $1/3$. The inner disk radius is taken to be the innermost stable circular orbit (ISCO) of a test particle, $r_{\rm in}=r_{\rm ISCO}=3r_{\rm s}$, where $r_{\rm s}=2G{\color[rgb]{0,0,0}M_{\rm PBH}}$ is the Schwarzschild radius. The outer radius is set to $r_{\rm out}=100r_{\rm s}$. The kinetic energy of a proton ($m_{p}=1.67\times 10^{-27}$ kg = 938.2 MeV) at a radius $r$ is determined based on the Keplerian velocity, $v(r)\propto 1/\sqrt{r}$, as
$E(r)\simeq\frac{1}{2}m_{p}v^{2}=782~{}{\rm keV}\left(\frac{r}{3r_{\rm s}}\right)^{-1}\left(\frac{f_{k}}{0.1}\right)^{2}.$ (16)
We take the velocity fraction of the outgoing proton wind to be $f_{k}=0.2$.
The energy distribution function of the outflow is given by [26]
@@
×$\times\left(\frac{n_{\rm H}}{1~{}{\rm cm^{-3}}}\right)\left(\frac{\tilde{v}}{10~{}{\rm km/s}}\right)^{-3}\left(\frac{r_{\rm out}}{r_{\rm s}}\right)^{-s}f_{v}^{2s}E^{-s-1},$ (17)
where the duty cycle factor is set to 1, $\tilde{v}=(v^{2}+c_{\rm s}^{2})^{1/2}\simeq 10$ km/s in Leo T (with $v$ and $c_{\rm s}\simeq 9$ km/s representing the PBH velocity relative to the gas and the temperature-dependent sound speed in the gas, respectively), and $f_{v}$ denotes the velocity distribution of the PBHs. [^4] Since the nature of the kinetic energy $E\propto v^{2}$ and the Keplerian velocity $v^{2}\propto 1/r$, $E\rightarrow 0$ corresponds to $r\rightarrow\infty$. In this study, $r_{\rm max}=r_{\rm out}=100r_{\rm s}\ll\infty$ and the energy distribution function $f(E)\propto 1/E^{s+1}$ cannot be divergent. Because the lack of significant relative bulk velocity in Leo T, we assume a standard Maxwell-Boltzmann distribution in polar velocity coordinates as
@@
where $\sigma_{v}$ is the velocity dispersion. Since the velocity dispersion of the PBHs is expected to match that of the gas, we take $\sigma_{v}=\sigma_{g}=6.9$ km/s, giving an average velocity $\bar{v}=\sigma_{v}\sqrt{8/\pi}\simeq 11.0$ km/s. The parameter $s$ has been estimated by simulations to range from 0.4 to 0.8. In this study, we use $s=0.7$.
We note that, since the radius $r$ in Eq. (16) is assumed to be $r=r_{\rm in}=3r_{\rm s}$ or $r=r_{\rm out}=100r_{\rm s}$, $E(r)$ is independent of the Schwarzschild radius $r_{\rm s}$. In a similar way, Eq.(17) is also independent of $r_{\rm s}$.
@@
It is important to note that heavy PBHs cannot constitute all of the DM. The maximum fraction of heavy PBH to DM is $f_{\rm PBH}\simeq 0.08$, which is only allowed within the narrow mass range of $1\leq M_{\rm PBH}/M_{\odot}\leq 30$ in green band in FIG. 2. In other mass regions, heavy PBHs can only make up a very small fraction of the DM [4, 26]. Therefore, similar to the case with light PBHs, the actual heating ratio in Leo T must be adjusted downward from what is shown in FIG. 2. However, unlike the light PBHs scenario, the case of heavy PBHs may yield significant results for the remainder of this paper, because $Q_{\rm PBH,outflow}$ can be comparable to $C$ in this mass range.
@@
2.5 Gas heating by dark photons
@@
In this subsection, we consider dark photons as the sole heating source for the gas in Leo T. Dark photons $A^{\prime}$ are kinematically mixed with the standard model photon $A$. The relevant Lagrangian for this study is [16]
$\mathcal{L}=-\frac{1}{4}{F}^{\mu\nu}F_{\mu\nu}-\frac{1}{4}{F^{\prime}}^{\mu\nu}F^{\prime}_{\mu\nu}+\frac{m_{A^{\prime}}^{2}}{2}A^{\prime\mu}A^{\prime}_{\mu}-\frac{g}{2}{F}^{\mu\nu}F^{\prime}_{\mu\nu}+J_{\rm em}^{\mu}A_{\mu},$ (20)
where $m_{A^{\prime}}$ denotes the mass of the dark photon, $g$ is the coupling parameter that describes the strength of the kinetic mixing, and $A_{\mu}$ ($A^{\prime}_{\mu}$) and $F_{\mu\nu}$ ($F^{\prime}_{\mu\nu}$) represent the gauge fields and field strengths of the visible (dark) photons, respectively. The electric current of the standard model is denoted by $J^{\mu}_{\rm em}$.
By redefining the gauge fields $A_{\mu}$ and $A^{\prime}_{\mu}$, we can remove the kinetic mixing term ${F}^{\mu\nu}F^{\prime}_{\mu\nu}$ [63, 15]. In these bases, the kinetic mixing translates into a direct coupling between the dark photon and the electric current. According to the formalism in Refs. [15, 28], the resulting Lagrangian is
$\mathcal{L}=-\frac{1}{4}{F}^{\mu\nu}F_{\mu\nu}-\frac{1}{4}{F^{\prime}}^{\mu\nu}F^{\prime}_{\mu\nu}+\frac{m_{A^{\prime}}^{2}}{2}A^{\prime\mu}A^{\prime}_{\mu}-\frac{e}{(1+g^{2})^{1/2}}J^{\mu}_{\rm em}\left(A_{\mu}+gA^{\prime}_{\mu}\right).$ (21)
@@
where $n_{e}$ denote the number density of electron, $T_{e}=6000$ K is the electron temperature, it is assumed to be virialized as in Eq.(18) which corresponds to $\sigma$ = 6.9 km/s, $\alpha_{\rm em}=1/137$ is the fine structure constant, and $\ln\Lambda_{\rm C}=0.5\ln[4\pi T_{e}^{3}/(\alpha_{\rm em}^{3}n_{e})]$ is the Coulomb logarithm, respectively.
@@
2.6 Gas heating by millicharged particles
In this subsection, we consider millicharged particles as the sole heating source for the gas in Leo T. In the millicharged DM model, a millicharged particle $\chi$ possesses a small electric charge $\epsilon e$ where $e$ is the electron charge [16]. There are two primary ways in which millicharged particles interact with the gas in Leo T [28].
The first way is the Coulomb-like interaction, involving the millicharged particles and free electrons or ions. In the gas of Leo T, almost all ions are ionized hydrogen atoms ${\rm H^{+}}$ (proton $p$). The total Rutherford scattering cross section for a millicharged particle $\chi$ interacting with an electron or ionized hydrogen is given by [28]
@@ 2.6 Gas heating by millicharged particles
$\lambda_{D}=\sqrt{\frac{{\color[rgb]{0,0,0}T_{e}}}{4\pi\alpha_{\rm em}\sum_{i}Z_{i}^{2}n_{i}}}.$ (26)
The second way to be available when the de Broglie wavelength of millicharged particle becomes smaller than the atom’s screening length. In this regime, the millicharged particle can interact with the nucleus of neutral atoms, as the nucleus is screened over a distance approximately equal to the Bohr radius $a_{0}$. For hydrogen atoms, this interaction becomes significant when $m_{\chi}\gtrsim 0.07$ GeV. In Leo T, neutral hydrogen and helium atoms are more prevalent than electrons or ions. Thus, for $m_{\chi}\gtrsim 0.07$ GeV, the dominant heat exchange by milicharged particles comes from neutral H and He atoms [28]. According to Refs. [64, 28], the total cross section for a millicharged particle $\chi$ interacting with hydrogen atom or helium atom is given by
$\sigma_{i\chi}=$ $\frac{2\pi Z_{i}^{2}\alpha_{\rm em}^{2}\epsilon^{2}}{\mu_{i}^{2}v_{\rm rel}^{4}}\left[\ln(1+4\mu_{i}^{2}v_{\rm rel}^{2}a^{2})-\frac{1}{1+(4\mu_{i}^{2}v_{\rm rel}^{2}a^{2})^{-1}}\right]\quad{\rm for}\ i=({\rm H,He}),$ (27)
@@
We assume that both the millicharged particle $\chi$ and the gas component $i$ follow Maxwell-Boltzmann distributions in Leo T. Under this assumption, the millicharged particle can heat (or cool) gas if $T_{i}\neq T_{\chi}$. Since there is no relative bulk velocity between the millicharged particle and gas in Leo T, the heating rate of the gas component $i$ per unit volume per unit time can be expressed as [28, 65]
@@
where $\sigma_{i\chi}^{(0)}=\sigma_{i\chi}/v_{\rm rel}^{n}$. According to $\sigma_{i\chi}\propto v^{-4}_{\rm rel}$ which is shown in Eq. (27), we use $n=-4$. Since the rms line-of-sight velocity dispersions of stars and the ${\rm H_{I}}$ gas are similar in Leo T, the velocity dispersions of the gas and the millicharged particle are approximately equal $v_{\chi}\simeq v_{\rm H_{I}}$. Consequently, thermal exchange between the millicharged particle and the gas depends on the mass differences because the millicharged particle and the proton, since $T_{\chi}-T_{\rm H_{I}}\simeq(m_{\chi}-m_{p})v_{\rm H_{I}}^{2}$. Thus, the heating rate will be small for $m_{\chi}\simeq m_{p}\simeq 1$ GeV.
@@
FIG. 4 illustrates this heating rate, assuming all DMs are composed of millicharged particles. The curve represents the heating rate under this assumption, while the horizontal line indicates the gas cooling rate.
3 Gas heating by both PBHs and DM particles
3.1 PBHs and dark photons
@@
The top panel of FIG. 5 illustrates the constraints on the PBH fraction $f_{\rm PBH}$ imposed by interstellar medium gas heating from both PBHs and dark photons in the Leo T dwarf galaxy. The black curve represents the upper limit of $f_{\rm PBH}$ when ignoring the gas heating effect of dark photons, as reported in Ref.[26]. The remaining three curves show the upper limit of $f_{\rm PBH}$ considering gas heating from both PBHs and dark photons. The red, blue, and green curves correspond to dark photon masses $m_{A^{\prime}}$ of $7\times 10^{-15}$ eV, $7.1\times 10^{-15}$ eV, and $7.2\times 10^{-15}$ eV, respectively. The excluded regions for $f_{\rm PBH}$ are indicated from gravitational waves [66] (GW : red), Icarus [67] (I : yellow), combined bounds from surviving astrophysical systems in Eridanus II [68], Segue 1 [69], and disruption of wide binaries [70] (S : orange), and Plank [71, 72] (P : gray).
@@
Next, we investigate the correlation between the upper bound on the mixing parameter $g$ and the mass of the dark photon $m_{A^{\prime}}$. We focus on the scenario where both PBHs and dark photons contribute effectively to the gas heating in Leo T. We consider PBH masses in the range $1\leq M_{\rm PBH}/M_{\odot}\leq 30$. In this mass range, the gas heating rate may be comparable to the cooling rate $C\simeq Q_{\rm PBH,outflow}$, see FIG. 2. Additionally, we use $f_{\rm PBH}=0.08$ as a benchmark case [4, 26].
The bottom panel of FIG. 5 shows the constraints on the kinetic mixing parameter $g$ from interstellar medium gas heating by PBHs and dark photons in the Leo T dwarf galaxy. The black curve represents the upper limit on $g$ when ignoring the gas heating effects of dark photons, as previously discussed in Ref. [28]. The remaining three curves depict the upper limit on $g$ considering gas heating by both PBHs and dark photons. The red, blue, and green curves correspond to PBH masses of $10M_{\odot}$, $20M_{\odot}$, and $30M_{\odot}$, respectively. Excluded regions for $g$ are also shown, including constraints from the Milky Way interstellar medium [15] (MW : orange), the gas cloud G33.4-8.0 [28] (G : yellow), and the CMB [63, 73] (CMB : gray).
We observe that most significant constraint on the kinetic mixing parameter is already obtained even when ignoring the gas heating effect of PBHs [28]. When considering the gas heating effect of PBHs in Leo T, this constraint is modified but remains the most significant.
3.2 PBHs and millicharged particles